# Could Planck Star Remnants be Dark Matter?

### v1

@@ 1 Introduction
One of the most profound unresolved questions in gravitational physics is the fate of matter undergoing complete gravitational collapse. In classical general relativity, such collapse inevitably leads to a singularity [1, 2, 3, 4, 5, 6, 7]. It is widely expected that quantum gravitational effects become important at the Planck density of $(c^{5}/hG^{2})\sim 10^{93}~{}{\rm g~{}cm^{-3}}$ and could resolve the singularity through new dynamics [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Loop Quantum Gravity (LQG) [18, 19, 20, 21, 22, 23], and in particular its application to cosmology in the form of Loop Quantum Cosmology (LQC) [24, 25, 26, 27, 28, 29, 30, 31], has provided an explicit manifestation of this expectation, replacing singularities with non-singular bounces for both cosmological and black hole contexts.
@@ 2 Quantum-Corrected Collapse
identifying the Schwarzschild areal radius $R$ at the boundary with the proper radius of the collapsing sphere. Our setup assumes that quantum gravitational effects are significant only in the high-density interior region [56, 57, 58] and become negligible outside the collapsing matter distribution. In particular, the Schwarzschild exterior solution remains a valid classical vacuum solution of General Relativity, since the quantum modifications in LQC are tied to regions of high curvature or Planck energy density. In this context, the vacuum exterior has a sub-Planckian curvature with a vanishing energy-momentum tensor and is not expected to receive significant corrections from quantum gravity effects. As such, matching a quantum-corrected FLRW interior to a classical Schwarzschild exterior is physically justified within the effective field theory approach  [^2] For details on how Effective Field Theory methods are employed, see [59, 60, 61]., where quantum gravity effects are localized and do not propagate beyond the matter distribution.
@@
The junction condition $[K_{ij}]=0$ implies that the difference in these quantities across $\Sigma$ must vanish and this leads to two coupled dynamical equations:
@@
Figure 1 shows the evolution of the collapsing sphere during the contraction phase governed by the FLRW metric under LQC. The radius of the collapsing sphere never reaches zero and the sphere begins to expand again after the bounce. This implies that the Schwarzschild exterior never collapses to a central singularity.
@@ 3 Planck Star Remnants as Dark Matter Candidates
In the bounce solution, Planck star remnants arise from non-singular quantum-corrected gravitational collapse. The Planck star scenario offers a compelling endpoint for the LQC-regulated dynamics studied earlier. The bounce halts the collapse at near-Planck densities, thus freezing the evolution and leaving behind a stable compact object. Once Hawking evaporation reduces the mass $M$ to the Planck mass, the bounce radius becomes of order the Schwarzschild radius and the long-lived relic could make dark matter candidates. One could ask as to why would Hawking evaporation stop and make a stable Planck-mass relic? In standard semiclassical theory Hawking evaporation proceeds down to arbitrarily small scales but in quantum gravity scenarios like ours, this behavior is expected to be modified. In this case what happens is that as the black hole approaches the Planck mass, quantum gravitational pressure can halt further evaporation replacing the classical singularity with a high-density core governed by LQG dynamics. This core, which would be the Planck star, has an interior governed by a bounce solution that prevents divergence of curvature invariants and terminates further collapse or explosive expansion. The outer geometry still satisfies Schwarzschild asymptotics but the quantum-corrected interior acts as a stable relic, consistent with the idea that Hawking evaporation ceases once the Planck regime is reached due to breakdown of semiclassical approximations.
@@
For relics formed in the very early universe, the dilution factor would be enormous. It would easily be exceeding $10^{64}$, which implies that the relics today would constitute cold dark matter. Moreover as the relic does not radiate or interact significantly after freeze-out, it behaves as pressureless matter and this ensures compatibility with cosmological observations. It is particularly in line with those constraining the cosmic microwave background and effective number of relativistic species $N_{\text{eff}}$ [69]. The entropy content of such relics is also likewise negligible compared to the thermal entropy of the cosmic microwave background or the cosmic neutrino background.
@@
In figure 2 we present a comparative plot of dark matter candidates in the mass–length scale plane. It is interesting to note that PSR occupy an intermediate region of this parameter space in the sense that they are significantly more massive than WIMPs or axions yet considerably lighter than primordial black holes, and their spatial distribution is neither highly clustered nor extremely rarefied. This balance positions PSR in a compelling ”sweet spot” with respect to mass and spatial abundance, which could potentially be allowing them to fulfill dark matter constraints without conflicting with structure formation or observational bounds.
We note that Planck mass relics have been discussed previously as potential dark matter candidates, most notably by MacGibbon [70] and follow-up work [71, 72, 73, 74, 75]. The previous models state that Hawking evaporation ceases once the mass reaches the Planck scale, leaving behind stable Planck mass relics. The motivation for such models is largely thermodynamic or based on semi-classical arguments for UV cutoffs without much said about the structure of these relics. In contrast, the PSR discussed in this work are compact objects with Schwarzschild-scale geometry and are derived from an explicit quantum gravitational scenario in LQG. The PSR arises when the classical singularity in gravitational collapse is resolved due to quantum geometry effects, resulting here in a bounce at the Planck density. The internal spacetime in our case is governed by an effective LQC-modified FLRW geometry, and the exterior is matched to a Schwarzschild metric. This is a fundamentally different picture than that envisioned by MacGibbon-type relics, which do not incorporate bounce dynamics or any non-trivial internal geometry.
@@ 4 Conclusions
The primary emphasis of earlier proposals of Planck Stars was on conceptual resolution of the singularity and information loss problems [34, 35]. These works also considered white hole remnants formed via black hole evaporation and even suggested that Planck-scale cores might tunnel into expanding white holes over long timescales [36, 37]. Our approach is substantially different as we implement tools from loop quantum theory directly into the collapsing interior geometry and consider its matching to a Schwarzschild exterior via Israel junction conditions. This framework allows us to demonstrate that a bounce occurs at Planckian density, halting collapse and there could be a formation of a non-singular PSR in this case. Moreover while others have considered a range of small masses [38, 39, 40, 69], our model considers a fixed Planck-scale mass for the PSR and this places the remnant in a well-defined region of parameter space with calculable number density and gravitational signatures.

### v2

@@ 1 Introduction
One of the most profound unresolved questions in gravitational physics is the fate of matter undergoing complete gravitational collapse. In classical general relativity, such collapse inevitably leads to a singularity [1, 2, 3, 4, 5, 6, 7]. It is widely expected that quantum gravitational effects become important at the Planck density of $(c^{5}/hG^{2})\sim 10^{93}~{}{\rm g~{}cm^{-3}}$ and could resolve the singularity through new dynamics [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Loop Quantum Gravity (LQG) [18, 19, 20, 21, 22, 23], and in particular its application to cosmology in the form of Loop Quantum Cosmology (LQC) [24, 25, 26, 27, 28, 29, 30, 31, 32], has provided an explicit manifestation of this expectation, replacing singularities with non-singular bounces for both cosmological and black hole contexts.
@@ 2 Quantum-Corrected Collapse
identifying the Schwarzschild areal radius $R$ at the boundary with the proper radius of the collapsing sphere. We also note that while $r_{0}$ is indeed constant and labels the comoving boundary of the FLRW region, the actual matching is done at a constant-time hypersurface across the bounce. Our setup assumes that quantum gravitational effects are significant only in the high-density interior region [57, 58, 59] and become negligible outside the collapsing matter distribution. In particular, the Schwarzschild exterior solution remains a valid classical vacuum solution of General Relativity, since the quantum modifications in LQC are tied to regions of high curvature or Planck energy density. In this context, the vacuum exterior has a sub-Planckian curvature with a vanishing energy-momentum tensor and is not expected to receive significant corrections from quantum gravity effects. As such, matching a quantum-corrected FLRW interior to a classical Schwarzschild exterior  [^2] While our analysis assumes a non-rotating, spherically sym.metric collapse,which is mostly fine for PBHs we note that extending this framework to include angular momentum remains a major challenge. The matching of a rotating interior to an exterior Kerr or cosmologically modified Kerr geometry is a nontrivial problem [60, 61]. is physically justified within the effective field theory approach  [^3] For details on how Effective Field Theory methods are employed, see [62, 63, 64]., where quantum gravity effects are localized and do not propagate beyond the matter distribution.
@@
We would also like to clarify about the nature of extrinsic curvatures here. Because our matching surface is spacelike (at fixed proper time in the interior), the relevant components of the extrinsic curvature are the spatial components (such as $K_{rr},K_{\theta\theta}$, etc.). The appearance of $K_{rr}$ is thus natural and necessary in our context and one would be right to think that if we were matching across a timelike worldtube then $K_{tt}$ and angular components would be the primary ones involved but here we are matching across a constant-time spatial slice, and hence the inclusion of $K_{rr}$ and the like is appropriate  [^4] For more details on these conditions, one can see [65, 66].. The junction condition $[K_{ij}]=0$ implies that the difference in these quantities across $\Sigma$ must vanish and this leads to two coupled dynamical equations:
@@
Figure 1 shows the evolution of the collapsing sphere during the contraction phase governed by the FLRW metric under LQC  [^5] The radius is normalized by its initial value $R(t=0)$, and time is expressed in Planck units $t/t_{\text{Pl}}$ so this normalization allows for dimensionless analysis of the collapse and bounce dynamics and allows one to scale both epochs properly in the figure.. The radius of the collapsing sphere never reaches zero and the sphere begins to expand again after the bounce. This implies that the Schwarzschild exterior never collapses to a central singularity.
@@ 3 Planck Star Remnants as Dark Matter Candidates
In the bounce solution, Planck star remnants arise from non-singular quantum-corrected gravitational collapse. The Planck star scenario offers a compelling endpoint for the LQC-regulated dynamics studied earlier. The bounce halts the collapse at near-Planck densities, thus freezing the evolution and leaving behind a stable compact object. Once Hawking evaporation reduces the mass $M$ to the Planck mass  [^8] We note that semiclassical arguments based on black hole entropy suggest that deviations from standard Hawking evaporation may begin well before the Planck mass is reached, most notably around the Page time, when half the initial entropy has been radiated. One such proposed modification is the recently advocated memory burden effect [73], which could significantly alter the late time dynamics and potentially prevent the formation of stable remnants. While our analysis focuses on the end stage dynamics near the Planck scale where loop quantum gravity effects dominate, we acknowledge that incorporating earlier quantum corrections such as those suggested by the memory burden mechanism is an important open direction for future work., the bounce radius becomes of order the Schwarzschild radius and the long-lived relic could make dark matter candidates. One could ask as to why would Hawking evaporation stop and make a stable Planck-mass relic? In standard semiclassical theory Hawking evaporation proceeds down to arbitrarily small scales but in quantum gravity scenarios like ours, this behavior is expected to be modified. In this case what happens is that as the black hole approaches the Planck mass, quantum gravitational pressure can halt further evaporation replacing the classical singularity with a high-density core governed by LQG dynamics. This core, which would be the Planck star, has an interior governed by a bounce solution that prevents divergence of curvature invariants and terminates further collapse or explosive expansion. The outer geometry still satisfies Schwarzschild asymptotics but the quantum-corrected interior acts as a stable relic, consistent with the idea that Hawking evaporation ceases once the Planck regime is reached due to breakdown of semiclassical approximations.
@@
For relics formed in the very early universe, the dilution factor would be enormous. It would easily be exceeding $10^{64}$, which implies that the relics today would constitute cold dark matter. To better justify the cold nature of PSR we note that while these relics are non-radiating and non-interacting after formation and their ”coldness” in a cosmological sense refers specifically to their negligible velocity dispersion and short free-streaming length. Any residual thermal energy from their evaporation era formation would be enormously redshifted by today with the suppression factor in (38), making their kinetic contribution to structure formation negligible and so this places them well within the cold dark matter regime. Moreover as the relic does not radiate or interact significantly after freeze-out, it behaves as pressureless matter and this ensures compatibility with cosmological observations. It is particularly in line with those constraining the cosmic microwave background and effective number of relativistic species $N_{\text{eff}}$ [75]. The entropy content of such relics is also likewise negligible compared to the thermal entropy of the cosmic microwave background or the cosmic neutrino background.
@@
In figure 2 we present a comparative plot of dark matter candidates in the mass–length scale plane. It is interesting to note that PSR occupy an intermediate region of this parameter space in the sense that they are significantly more massive than WIMPs or axions  [^9] In this figure we have plotted a representative axion mass of order $\sim 10^{-5}$eV which is consistent with conventional QCD axion models but we do acknowledge that axion-like particles in string inspired scenarios or other extensions could span a vastly broader mass range, from $10^{-22}$eV up to the MeV scale [76, 77, 78]. yet considerably lighter than primordial black holes, and their spatial distribution is neither highly clustered nor extremely rarefied. This balance positions PSR in a compelling ”sweet spot” with respect to mass and spatial abundance, which could potentially be allowing them to fulfill dark matter constraints without conflicting with structure formation or observational bounds.
We note that Planck mass relics have been discussed previously as potential dark matter candidates, most notably by MacGibbon [79] and follow-up work [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90] and the consideration of remnant type dark matter was recently seen in [91]. The previous models state that Hawking evaporation ceases once the mass reaches the Planck scale, leaving behind stable Planck mass relics. The motivation for such models is largely thermodynamic or based on semi-classical arguments for UV cutoffs without much said about the structure of these relics. In contrast, the PSR discussed in this work are compact objects with Schwarzschild-scale geometry and are derived from an explicit quantum gravitational scenario in LQG. The PSR arises when the classical singularity in gravitational collapse is resolved due to quantum geometry effects, resulting here in a bounce at the Planck density. The internal spacetime in our case is governed by an effective LQC-modified FLRW geometry, and the exterior is matched to a Schwarzschild metric. This is a fundamentally different picture than that envisioned by MacGibbon-type relics, which do not incorporate bounce dynamics or any non-trivial internal geometry.
@@ 4 Conclusions
The primary emphasis of earlier proposals of Planck Stars was on conceptual resolution of the singularity and information loss problems [35, 36]. These works also considered white hole remnants formed via black hole evaporation and even suggested that Planck-scale cores might tunnel into expanding white holes over long timescales [37, 38]. Our approach is substantially different as we implement tools from loop quantum theory directly into the collapsing interior geometry and consider its matching to a Schwarzschild exterior via Israel junction conditions. This framework allows us to demonstrate that a bounce occurs at Planckian density, halting collapse and there could be a formation of a non-singular PSR in this case. Moreover while others have considered a range of small masses [39, 40, 41, 75], our model considers a fixed Planck-scale mass for the PSR and this places the remnant in a well-defined region of parameter space with calculable number density and gravitational signatures. A viable question which may also arise could be about the observational features of PSR and whether one can expect such features to even exist. In this regard, we would like to remark that a follow up work of ours on this subject addresses this aspect aptly as well.
We would also like to note that in our scenario the key idea is that PSR is not treated as a generic or structureless endpoint, but as the result of a well defined dynamical bounce described by effective Loop quantum dynamics. The internal structure of the remnant encodes a non trivial geometric history and to be specific, it lies in the bounce evolution of the collapsing matter. In this sense the PSR is not a degenerate endpoint of many initial conditions but a unique outcome of a unitary evolution governed by quantum gravitational corrections. The entropy that accumulates in the Hawking radiation reflects entanglement with the interior degrees of freedom of the remnant, which are governed by a high-density quantum geometry  [^10] For some details of the entropy structure for Hawking radiation, see [92, 93].. While we do not yet have a complete microscopic accounting of the remnant’s entropy, the LQC framework in general implies that the bounce retains information about the initial state and this opens up the possibility that the fine-grained entropy of the remnant need not be small even though its geometric entropy (in the Bekenstein sense) is. Our picture would thus be consistent with a pure global state when both radiation and interior remnant are considered together. It may even be an interesting endeavor to pursue a different work which is focused on the details of the entropy structure of PSR, particularly treating its entropy in LQG consistent ways.